OpenStudy (anonymous):

4to the ? power times 4 to the third power = 4 to the 12 power what is the question marks power?

OpenStudy (anonymous):

PLease Help!

OpenStudy (anonymous):

\[4^x*4^3=4^{12}\]mult like bases add exponents\[4^{x+3}=4^{12}\]equate bases by a theorem\[x+3=12\]subtract 3\[x=9\]this is the missing exponent

OpenStudy (anonymous):

Check the result\[4^9*4^3=4^{9+3}=4^{12}\]

OpenStudy (anonymous):

the theorem referenced above is \[b^x=b^y \iff x=y\]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

my pleasure :})